A manufacturer claims that at least 95% of the equipment that she supplied to a
factory confirmed to the specification. An examination of a sample of 200 pieces
of equipment revealed 18 were faulty. Test for her claim at levels of significance (a)
0.01 and (b) 0.05

Answer by stanbon(75887)

A manufacturer claims that at least 95% of the equipment that she supplied to a
factory conformed to the specification. An examination of a sample of 200 pieces
of equipment revealed 18 were faulty. Test for her claim at levels of significance
(a) 0.01
Ho: p >= 0.95 (claim)
Ha: p < 0.95
-----
p-hat = 182/200 = 0.91
z(0.41) = (0.91-0.95)/sqrt[0.95*0.05/200] = -0.04/sqrt[0.95*0.05/200]
= -2.5955
--------------
p-value = P(z < -2.5955) = 0.0047
------
Conclusion: Since the p-value is less than 1%, reject Ho ; reject the claim.

(b) 0.05
Conclusion:: Since the p-value is less than 5%, reject Ho ; reject the claim.
Cheers,
Stan H.